OpenStudy (ineedhelpnow):

solve for H 80=(1/2)(16)(H)

OpenStudy (saifoo.khan):

cancel 16 and 2. you will get 8. 80 = 1*8 *H 80 = 8H \[\frac{80}{8}=H\] \[h = 10\]